OpenStudy (spicytaco1):

What are the square roots of 256? A. 15 and –15 B. 16 and –16 C. 25 and –25 D. 128 and –128

OpenStudy (pugtato):

16x16= 256 so the answer is most likely B

OpenStudy (decentnabeel):

\[\sqrt{256}\] \[\mathrm{Factor\:the\:number:\:}\:256=16^2\] \[=\sqrt{16^2}\] \[\mathrm{Apply\:radical\:rule:\:}\sqrt[n]{a^n}=a\] \[\sqrt{16^2}=16\]

OpenStudy (decentnabeel):

@Spicytaco1 are you gotit

OpenStudy (spicytaco1):

Thanks :D

OpenStudy (decentnabeel):

Welcome :)